# Birational complexity and conic fibrations

### v1

@@ 5.6. Weil index of Fano varieties
The Weil index $i(X)$ of a Fano variety $X$ is the largest positive integer for which we can write $-K_{X}\sim iW$ with $W$ a Weil divisor. If $X$ is smooth, then $i(X)\leq n+1$ and $X\simeq\mathbb{P}^{n}$ whenever $i(X)>n$ (see, e.g., [24]). However, if $X$ is singular, the Weil index can be arbitrarily large. For instance $\mathbb{P}(1,1,n)$ has Weil index $n$. Thus, a priori, the Weil index can be large and not reflect in the geometry of the Fano variety. The following theorem shows that $n$-dimensional Fano varieties with large Weil index (compared to $n$) have non-maximal alteration conic complexity.
@@
[…]
By construction, the degree of each $f_{i}$ is positive. We write $g_{i}:=f_{1}\circ\dots\circ f_{i}$. Thus, we have $\deg(g_{i})>\deg(g_{i-1})$ for each $i$. Let $h_{i}\colon Y_{i}\rightarrow X$ be the Galois closure of $g_{i}$. Then, we have $\deg(h_{i})\geq\deg(g_{i})$. Furthermore, if $(Y_{i},\Delta_{i})$ is the log pull-back of $(X,B)$ to $Y_{i}$, then $h_{i}\colon(Y_{i},\Delta_{i})\rightarrow(X,B)$ is a crepant finite Galois morphism. By Lemma 2.17, $Y_{i}$ is of Fano type and $\Delta_{i}$ is a $1$-complement. There is a finite group $G_{i}$ with $|G_{i}|\geq 2^{i}$ acting on $(Y_{i},\Delta_{i})$. By [8, Theorem 3], there is a normal abelian subgroup $A_{i}\leqslant G_{i}$ of rank at most $n$ and index at most $k_{0}(n)$. Let $m_{0}:=m_{0}(n,\{0,1\})$ be as in Theorem 5.6. If $\deg(g_{i})>k_{0}m_{0}^{n}$, then $(Y_{i},\Delta_{i})$ admits the action of $\mathbb{Z}_{l}$ with $l>m_{0}$. Thus, by Theorem 5.6, we conclude that $k_{\rm bir}(Y_{i},\Delta_{i})<n$ so $k_{\rm alt}(X,B)<n$. We conclude that the sequence of finite covers stops at $(X_{s},B_{s})$ for some $s$. Furthermore, the degree of the finite morphism $X_{s}\rightarrow X$, which we may assume Galois, is bounded above by $k_{0}m_{0}^{n}$. By construction, the components of $B_{s}$ generate the lattice $\langle B_{s}\rangle\cap{\rm Cl}(X_{s})$. Thus, it suffices to set $(Y,B_{Y}):=(X_{s},B_{s})$. ∎
@@ 5.7. Complexity of klt singularities
Let $(X;x)$ be a $n$-dimensional $\mathbb{Q}$-factorial klt singularity. Let $B$ be a $1$-complement of the singularity. Let $b$ be the number of components of $B$ through $x\in X$. Let $\nu$ be a log canonical place with $c_{X}(\nu)=x$. By [32, Theorem 1], there exists a projective birational morphism $\pi\colon Y\rightarrow X$ from a normal $\mathbb{Q}$-factorial variety which is an isomorphism over $X\setminus\{x\}$ and $\phi^{-1}(x)=E_{\nu}$ is a unique prime divisor corresponding to $\nu$. Let $\pi^{*}(K_{X}+B)=K_{Y}+B_{Y}+E_{\nu}$. The divisor $E_{\nu}$ may not be normal. However, we know that $E_{\nu}$ is an slc variety. Hence, there exists a projective birational morphism $\phi\colon Z\rightarrow X$ satisfying the following conditions:
@@ 6. Examples and Questions
[…]
@@
is log Calabi–Yau and its complexity is $d-1$. So, the birational complexity of $X_{d}$ is at most $d-1$. By Theorem 1.1, the variety $X_{d}$ has a birational model $X^{\prime}_{d}$ that admits a sequence of Mori fiber spaces
[…]
of which at least $n-d+1$ are conic fibrations. On the other hand, Kollár proved that whenever $X_{d}$ is very general and $d\geq 3\lceil\frac{n+3}{4}\rceil$ the variety $X_{d}$ does not admit a birational conic fibration (see [25] and [26, Theorem V.5.14]). Thus, the first Mori fiber space $\phi_{1}$ is not a conic fibration. In the case of cubic $3$-folds $X_{3}\subset\mathbb{P}^{4}$ our result simply recovers the birational pencil structure described by Pukhlikov [39]. In this case, the conic fibration is the structure morphism of the base of the pencil.
@@
Finally, Theorem 1.1 states that a log Calabi–Yau pair $(X,B)$ of dimension $n$ and birational complexity $c$ has a birational model that admits a tower of Mori fiber spaces of which at least $n-c$ are conic fibrations. In the case of smooth Fano hypersurfaces, we can make sure that the conic fibrations are precisely the last $n-c$ Mori fiber spaces. This seems to be the optimal situation. Indeed, in this case, the base of the fibrations is a Bott $\mathbb{Q}$-tower (see Theorem 1.5). It is not clear whether this is the case for other log Calabi–Yau pairs.
@@
Let $(X,B)$ be a log Calabi–Yau pair of dimension $n$ and birational complexity $c$. What are the possible positions of the $n-c$ conic fibrations in birational towers of Mori fiber spaces for $(X,B)$?

### v2

@@ 5.7. Weil index of Fano varieties
The Weil index $i(X)$ of a Fano variety $X$ is the largest positive integer for which we can write $K_{X}\sim iW$ with $W$ a Weil divisor. If $X$ is smooth, then $i(X)\leq n+1$ and $X\simeq\mathbb{P}^{n}$ whenever $i(X)>n$ (see, e.g., [24]). However, if $X$ is singular, the Weil index can be arbitrarily large. For instance $\mathbb{P}(1,1,n)$ has Weil index $n$. Thus, a priori, the Weil index can be large and not reflect in the geometry of the Fano variety. The following theorem shows that $n$-dimensional Fano varieties with large Weil index (compared to $n$) have non-maximal alteration conic complexity.
@@
${(X,B)}$${(X_{1},B_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{1}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{2}}$${(X_{r},B_{r})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{r}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{r+1}}$
By construction, the degree of each $f_{i}$ is positive. We write $g_{i}:=f_{1}\circ\dots\circ f_{i}$. Thus, we have $\deg(g_{i})>\deg(g_{i-1})$ for each $i$. Let $h_{i}\colon Y_{i}\rightarrow X$ be the Galois closure of $g_{i}$. Then, we have $\deg(h_{i})\geq\deg(g_{i})$. Furthermore, if $(Y_{i},\Delta_{i})$ is the log pull-back of $(X,B)$ to $Y_{i}$, then $h_{i}\colon(Y_{i},\Delta_{i})\rightarrow(X,B)$ is a crepant finite Galois morphism. By Lemma 2.17, $Y_{i}$ is of Fano type and $\Delta_{i}$ is a $1$-complement. There is a finite group $G_{i}$ with $|G_{i}|\geq 2^{i}$ acting on $(Y_{i},\Delta_{i})$. By [7, Theorem 3], there is a normal abelian subgroup $A_{i}\leqslant G_{i}$ of rank at most $n$ and index at most $k_{0}(n)$. Let $m_{0}:=m_{0}(n,\{0,1\})$ be as in Theorem 5.6. If $\deg(g_{i})>k_{0}m_{0}^{n}$, then $(Y_{i},\Delta_{i})$ admits the action of $\mathbb{Z}_{l}$ with $l>m_{0}$. Thus, by Theorem 5.6, we conclude that $k_{\rm bir}(Y_{i},\Delta_{i})<n$ so $k_{\rm alt}(X,B)<n$. We conclude that the sequence of finite covers stops at $(X_{s},B_{s})$ for some $s$. Furthermore, the degree of the finite morphism $X_{s}\rightarrow X$, which we may assume Galois, is bounded above by $k_{0}m_{0}^{n}$. By construction, the components of $B_{s}$ generate the lattice $\langle B_{s}\rangle\cap{\rm Cl}(X_{s})$. Thus, it suffices to set $(Y,B_{Y}):=(X_{s},B_{s})$. ∎
@@ 5.8. Complexity of klt singularities
Let $(X;x)$ be a $n$-dimensional $\mathbb{Q}$-factorial klt singularity. Let $B$ be a $1$-complement of the singularity. Let $b$ be the number of components of $B$ through $x\in X$. Let $\nu$ be a log canonical place with $c_{X}(\nu)=x$. By [34, Theorem 1], there exists a projective birational morphism $\pi\colon Y\rightarrow X$ from a normal $\mathbb{Q}$-factorial variety which is an isomorphism over $X\setminus\{x\}$ and $\phi^{-1}(x)=E_{\nu}$ is a unique prime divisor corresponding to $\nu$. Let $\pi^{*}(K_{X}+B)=K_{Y}+B_{Y}+E_{\nu}$. The divisor $E_{\nu}$ may not be normal. However, we know that $E_{\nu}$ is an slc variety. Hence, there exists a projective birational morphism $\phi\colon Z\rightarrow X$ satisfying the following conditions:
@@ 6. Examples and Questions
Therefore, $\mathbf{M}$ is a torsion b-nef divisor provided $m>2\lambda_{0}^{-1}$. Otherwise from (6.2), we conclude $\deg\left(\mathbf{M_{\mathbb{P}^{1}}}\right)>2$ leading to a contradiction. On the other hand, if $m>2\lambda_{0}^{-1}$, then $B$ is supported on $\{0\}$ and $\{\infty\}$, as the coefficients of $B$ are in $\Omega$. Thus, if $m\geq\min\{7,2\lambda_{0}^{-1}\}$, then $\mathbf{M}$ is a torsion b-nef divisor and $B$ is supported on $\{0\}+\{\infty\}$ which implies the isomorphism (6.1).
@@
is log Calabi–Yau and its complexity is d1d-1. So, the birational complexity of $X_{d}$ is at most d1d-1. By Theorem 1.2, the variety $X_{d}$ has a birational model $X^{\prime}_{d}$ that admits a sequence of Mori fiber spaces
${X^{\prime}_{d}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${X^{\prime}_{d,1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{2}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{k}}$${X^{\prime}_{d,k}}$
of which at least nd+1n-d+1 are conic fibrations. On the other hand, Kollár proved that whenever $X_{d}$ is very general and $d\geq 3\lceil\frac{n+3}{4}\rceil$ the variety $X_{d}$ does not admit a birational conic fibration (see [27] and [28, Theorem V.5.14]). Thus, the first Mori fiber space $\phi_{1}$ is not a conic fibration. In the case of cubic $3$-folds $X_{3}\subset\mathbb{P}^{4}$ our result simply recovers the birational pencil structure described by Pukhlikov [41]. In this case, the conic fibration is the structure morphism of the base of the pencil.
@@
Finally, Theorem 1.2 states that a log Calabi–Yau pair $(X,B)$ of dimension $n$ and birational complexity $c$ has a birational model that admits a tower of Mori fiber spaces of which at least ncn-c are conic fibrations. In the case of smooth Fano hypersurfaces, we can make sure that the conic fibrations are precisely the last ncn-c Mori fiber spaces. This seems to be the optimal situation. Indeed, in this case, the base of the fibrations is a Bott $\mathbb{Q}$-tower (see Theorem 1.6). It is not clear whether this is the case for other log Calabi–Yau pairs.
@@
Let $(X,B)$ be a log Calabi–Yau pair of dimension $n$ and birational complexity $c$. What are the possible positions of the ncn-c conic fibrations in birational towers of Mori fiber spaces for $(X,B)$?